STUDIES IN LOGIC, GRAMMAR AND RHETORIC 1 (14)

Jerzy Kopania

DESCARTES’ GREAT THESIS ON NATURE

I

Carrying out the reasoning process of total methodical doubt Descartes came to the conclusion that the only primary and undoubtful thesis was: “I think, therefore I am”. However, what can I think of if everything is doubted with respect to its existence? It is clear that I can think of my own thinking only. Thinking is thus a totally immanent process, which determines its self-sufficiency, because directing it toward external reality would be incompatible with immanence. In other words: the inner content of the mind is the only object of cognition.

But if thinking is a totally immanent process, is it possible to demonstrate that there exists anything except the mind of knower? That is how a certain problem arose in Descartes’ system, a problem which had never been posed before – at least not so clearly and consciously. Descartes formulated the question of the existence of the universe and it was a philosophical question, that is to say: a theoretical, not a practical one.

Contrary to the scholastic followers of Aristotle, which one and all based on the uncritical confidence in the objective obviousness of the world, Descartes was of the opinion that philosophy should prove the existence of the independent world rather than assume it dogmatically. In order to produce such a proof it was necessary to assume that the reality of the mind was not augmented by anything derived from the reality of the world, which meant assuming that the human mind had a complete set of innate ideas at its disposal, ideas being both the only object of cognition and the necessary condition of its possibility and validity. Lack of this assumption portends a vicious circle in demonstration: by claiming that the content of the mind is augmented by some elements of extra-mental reality one assumes the existence of this reality, which is exactly what was to be proved.

Jerzy Kopania

Additional confirmation of such a standpoint was the mechanistic vision of the universe held by Descartes. If all events taking place in nature are of purely mechanistic character, than anything that is not mechanistic cannot affect the mind from outside. And since the mind is aware of not only this kind of influence – we experience the objects, and not the stimuli, as Kant will say later – everything else must be innate in it. In other words: the elements of independent world only create an opportunity for the mind to become aware of what is innate in it. In this way the mechanistic standpoint started to play the role of a verifying factor of the idealistic viewpoint in philosophy.

II

It might seem that some basic contradiction emerged in Descartes’ reasoning: as philosopher he propagates the thesis that the only direct object of cognition is the inner content of the mind, while as man of science he does empirical research in natural reality in a variety of its aspects. If all cognitive operations are carried out on ideas only, what can the role of empirical research be?

Descartes’ great thesis on nature

Descartes thought that on the ground of his philosophy he proved the existence of the universe. Yet, the approach to it was exclusively indirect through the inner cognitive structures of the mind. That, however, brings about the appearance of some essential problem – how to distinguish that which is only possible (viz. only inside the mind) from that which is realized in the world outside the mind. The adequacy between the cognitive structure of the mind and the structure of the universe is not simply established but still has to be achieved. The point is that only one of the possible explanations is actually connected in a necessary way with the phenomenon being explained and it has to be determined somehow which one it is. An ordinary experiment is not sufficient here and one has to resort to crucial experiments. That is why emphasizing in Discourse on Method the fact that he does not know any physical phenomena which could not be explained by means of his principles, at the same time Descartes expresses his reservation: “But I must also confess that the power of nature is so ample and so vast, and these principles are so simple and general that I observed hardly any particular effect as to which I could not at once recognize that it might be deduced from the principles in many different ways; and my greatest difficulty is usually to discover in which of these ways the effect does depend upon them. As to that, I do not know any other plan but again to try to find experiments of such a nature that their result is not the same if it has to be explained by one of the methods, as it would be if explained by the other.”^{1} Thus the point is in planning experiments in such a way that they might produce various results, depending on the methods of explanation employed. The difficulty of which Descartes speaks consists in time consumption and big costs of carrying out such experiments. Hence, it is necessary to limit them to several selected issues and to cooperate with other researchers.

The investigative method of Descartes was in its assumption supposed to be a synthesis of deduction and experiment because it was to discover the necessary and unchangeable relation between the datum of experiment and the quaesitum of theory, which seen from the methodological viewpoint resolved itself into discovering the essential connection of foreseeing and explaining. On the basis of experimental data subjected to initial analysis and classification, hypothetical ,suppositions. concerning their cause are introduced. The complexity and magnitude of nature cause the suppositions to be varied and numerous, and therefore empirical conclusions deduced from them will also be varied and numerous. Hence, to eliminate false suppositions one needs to use the crucial experiment (Bacon’s instantia crucis) and – speaking in Mill’s language – the methods of agreement, difference, and concomittant changes. If the elements of a theory precisely correspond with the elements of a phenomenon, then the true case was discovered. In such a way the theory was on the one hand supposed to explain facts, while on the other – facts were to confirm the theory.

III

The question of how to understand the proposition that theory explains facts, being at the same time confirmed by them, Descartes touched upon in the Discourse on Method. “If some of the matters of which I spoke in the beginning of the Dioptrics and Meteors should at first sight give offence because I call them hypotheses and do not appear to care about their proof, let them have the patience to read these in entirety, and I hope that they will find themselves satisfied. For it appears to me that the reasonings are so mutually interwoven, that as the later ones are demonstrated by the earlier,

1 AT VI, 64-65; HR I, 121. References beginning with AT are by volume and page to Oeuvres de Descartes, ed. by Adam and Tannery, Paris 1974-1983, and with HR are to The Philosophical Works of Descartes, transl. by E. S. Haldane and G. R. T. Rose, Cambridge University Press, 1967.

which are their causes, the earlier are reciprocally demonstrated by the later which are their effects. And it must not be imagined that in this I commit the fallacy which logicians name arguing in a circle, for, since experience renders the greater part of these effects very certain, the causes from which I deduce them do not so much serve to prove their existence as to explain them; on the other hand, the causes are explained by the effects.”^{2}

The above text could cause misunderstanding. Jean-Baptiste Morin had his doubts in this respect and he expressed them (along with a number of other scientific problems) in a letter to Descartes written on February 22, 1638. In his reply Descartes tries togive explanation of the terms used in the above quotation from the Discourse. He naturally admits that proving effects by means of causes and then proving causes by the same effects actually means a vicious circle in reasoning. However, in the case under consideration, the problem consists in something else. “I do not agree that it is circular to explain effects by a cause and then prove the cause by the effects; because there is a big difference between proving and explaining. I should add that the word ,demonstrate. can be used to signify either, if it is used according to common usage and not in the technical philosophical sense. I should add also that there is nothing circular in proving a cause by several effects which are independently known, and then proving certain other effects from their cause.”^{3}

Descartes defends himself pointing out that he used the word “demonstrate” (démontrer) in its everyday and not technical sense; whenever he means demonstration in mathematical or philosophical sense, he uses the form “prove” (prouver) and it is this understanding of demonstration that he contrasts with explanation. Stating in the Discourse that effects demonstrates causes and causes demnostrate effects, Descartes includes two meanings (viz. prove and explain) in one word “demonstrate”. In spite of that he does not think he might be charged with ambiguous pronouncements because of the immediately expressed reservation that the causes from which effects are deduced serve not so much as a means of proving those effects, but rather as a means of explaining them. Hence it is the causes that are proved by effects. Descartes says in addition: “And I put ,serve not so much to prove them. rather then ,do not serve at all., so that people could tell that each of these effects could also be proved from other effects. I do not see what other term I could have used to explain

2 AT VI, 76; HR I, 128-129.

3 AT II, 198; K, 57-58. References beginning with K are by page to: Descartes, Philosophical Letters, transl. and ed. by A. Kenny, Oxford, Clarendon Press, 1970.

myself better.”^{4}

The explanations given by Descartes are formulated in a somewhat complicated stylistics, though one can grasp his understanding of the relations between cause and effect. Between statement C ascertaining a cause and statement E ascertaining an effect there must be the relation of logical inference. Then statement C is the reason of statement E and statement E is the consequence of statement C, viz. statement E follows from statement C. That is how one should understand Descartes’ ascertainment that he deduces effects from causes. The relation of inference can be characterized only syntactically, though semantic characteristics can also be given. On the grounds of pure syntax one determines inference, the so-called syntactic derivability, in the following way: statement S1 is a logical consequence of statement S2 if and only if the implication whose antecedent is S1 and consequent is S2 is a substitution of a logical tautology, viz. of such a sentential schema which turns into a true statement at each substitution. Semantic consequence, in turn, is characterized by means of the notion of a model: statement S1 is a semantic consequence of statement S2 if and only if in each model in which S1 is true S2 is also true. Syntactic consequence is thus characterized wit respect to the given deductive system while semantic consequence with respect to the models of that system. When Descartes says that he deduces an effect from a cause, such a pronouncement should be interpreted as referring to semantic consequence. After all, he does not build a formal system on the ground of which inference would take place. He has at his disposal a set of laws composed of the so-called eternal truths.^{5} It is a heterogenous set made of logical and mathematical truths as well as ontological theses.If, however, we treat this set as a specific system, the world of nature will be its model. Then the ascertainment that statement E follows from statement C should be understood in the sense that statement C cannot be true and statement E false in any conceivable world being a model for those laws. This is of course secured by logical relations between the elements of the set of truths and it is in this sense that semantic consequence would be equivalent to syntactic consequence if we treated that set as a formal system.

Then one may speak of deducing, just as Descartes does. Here we have a situation in which only true statements follow from true statements and

4 AT II, 198; K, 58.

5 These laws are eternal but created by God. Descartes identified the creation of laws with the creation of essences, viz. the eternal truth is the same as the essence of what the truth is about, e.g. the essence of the circle lies in its being a space enclosed by a curved line every point of which is equidistant from the center, and this thesis is the eternal truth about the circle.

thus our inference is infaliable, that is deductive. It is in this context that investigative procedure should be understood as consisting in selecting for a given effect such a cause from which it will logically follow. Of course, such a procedure is not any demonstration, and Descartes emphasizes it. The cause selected in such a way does not prove the effect. Proving refers to statements not recognized yet as true statements; they will be done so only after producing the proof. In other words: on the basis of statements recognized as true, the truthfulness of statements being their logical consequence is proved. In the case under consideration it is the statement ascertaining the effect that is recognized as true and not the statement ascertaining the cause. If statement C were in some way proved, thus at the same time recognized as true, and since E is a logical consequence of C, then statement E could be proved by means of statement C. That is the situation to which the latter of the quotations from Descartes refers. He emphasizes the fact that such a procedure would be recommended if any doubts concerning the truthfulness of the statement E arose. Demonstration is necessary when the truthfulness of the statement being demonstrated is not obvious.

A question arises immediately as to the purpose of selecting a cause for a known effect, viz. selecting a statement ascertaining the cause for the statement ascertaining the effect which is known to be true statement. The purpose is to give explanation of the effect; for even if we recognize the statement ascertaining the effect as true, we do not know the reason underlying the truthfulness. The explanations of some state of affairs consists in answering the question why this state of affairs came into being, which is the question concerning its cause. The question can only be answered in such a way that we give the statement whose logical consequence is the statement ascertaining that state of affairs. The direction of explanation is thus in accord with that of logical inference, i.e. deduction, and the explanation concerns the statement already recognized as true.

However, the very occurence of the relation of logical inference does not suffice, viz. it is not always so that if statement S2 is a consequent of statement S1 we may recognize S1 as explaining the state of affairs ascertained in S2. The occurence of the relation of logical inference between statements is a necessary condition, though not sufficient for the relation of explanation to occur between them. Moreover, the very explanatory statement must be of a special type, viz. it must possess a certain feature thanks to which it explains the statement that follows from it. In the case of one single fact the statement explaining it will be the statement acsertaining its cause. When we deal with a recurrent phenomenon, the statement explaining this phenomenon must be a general statement expressing some law of nature for it is only then that an adequate conditional statement will describe the relation of causal determination. In other words: statement C explains statement E if and only if statement C is a law whose logical consequence is statement E. The conditional statement “if C, then E” describes the relation of causal determination occurring between the state of affairs ascertained by the law expressed in statement C and the state of affairs ascertained by statement E. Obviously, because of the complex character of natural phenomena and the entaglement of explanatory statements in the explanatory theory, in the actual process of explanation we mostly deal not with a single statement C but with a conjunction of statements from which statement E follows.

The above settlements make it possible to give a certain interpretation of those parts of the Discourse and the analysed letter in which Descartes claims that the effect proves cause. The statement ascertaining the cause is selected in such a way that the statement ascertaining the effect might follow from it and an adequate conditional statement might be the description of the relation of causal determination thanks to which the explanation of the analysed state of affairs is given. How should one understand in this context the statement claiming that the effect proves the cause?

Proving consists in the fact that on the basis of statements recognized as true one demonstrates the truthfulness of statements which are their logical consequence. Hence, statement E ascertaining the effect is obviously recognized as true, but statement C ascertaining the cause does not follow from it (the direction of inference is opposite). If, then, the effect is to demonstrate the cause, the very statement E cannot be a premise; most likely the point is in some conjunction of statements where statement E is one of its elements and statement C is its logical consequence. Since the conditional statement “if C, then E” expresses the relation of causal determination, we may assume that an additional conjectural premise also concerns this relation.

Descartes was very firm in the way he propounded his thesis according to which no thing exists without the cause of its existence. He put forward the following thesis as his axiom: “Nothing exists concerning which the question may not be raised – ,What is the cause of its existence?”^{6} In the terminology derived from Aristotle the cause bringing about the realisation of being (the fact of beginning to be) used to be called the efficient cause. The acting of the efficient cause was expressed

6 AT VII, 164; HR II, 55.

in the form of the so-called principle of causality. Most generally it can be formulated sa follows: “everything that comes into being has its cause”, which is tantamount to saying: “every effect has its cause”. In the methodological procedure described by Descartes for statement E ascertaining the occurence of a certain state of affairs one selects reason C explaining this state. Consequently, one formulates a certain description of causal determination dependence. Thus, one assumes here the principle of causality, though without expressing it overtly, and after accepting statement C as explanatory we actually obtain two statements: (1) “the state of affairs ascertained in statement E did occur”, and (2) “if the state of affairs ascertained in statement E occurred, then its cause described in statement C also existed”. By virtue of the rule of ponendo ponens a logical consequence of the above two statements is the statement: (3) “as described in statement C, the cause of the state of affairs ascertained in statement E did exist”. Thus, the above inference is an act of proving because on the basis of statements (1) and (2) recognized as true one comes to recognize statement (3) as true, too. If Descartes claims that the effect proves the cause, then according to the presented interpretation one should understand it in such a way that the statement ascertaining the occurence of a given state of affairs proves the statement ascertaining the existence of the cause of this state of affairs.

IV

Any deductive inference is infaliable, but the infaliability concerns the course of inference and cannot guarantee the truthfulness of premises. Deductive inference makes it certain that the conclusion is a true statement if all premises are true statements. The problem, though, is that a true conclusion might be drawn from false premises. That is why the crucial experiment was of such vital importance: it enabled the elimination of all causes except one from the set of all those that were possible. A true conclusion could be inferred from a true premise then. It is striking that Descartes, who emphasized the heuristic futility of syllogistic (i.e. deduction), at the same time put so much attention to the deductive method of pursuing science, the method which scholastics belived to be the only permissible one. His standpoint becomes understandable when we take into consideration the difference in points of departure. If a conclusion obtained deductively contains in its substance only that which was already contained in the premises, the selection of premises becomes the decisive issue. Descartes’ premises were basically different from those of scholastics. Their point of departure was the external world in the way it appeared to them in their objectively obvious sensory perception. What Descartes started from was the inner reality of the mind. He was convinced that that was the difference which determined the efficiency of his investigative method.

It is true that it is easy to select many separate causes for many effects – Descartes wrote to Morin. However, when one is not concerned with seeking for a direct cause of a single state of affairs, it is not easy to select the one and only cause for many effects. In such cases one needs to select the one cause from which all effects can be deduced and at the same time it is proved to be the true cause of them all. Then Descartes emphasizes the fact that all the causes considered by him are of such a kind. On the grounds of physics efforts had always been made to envisage some causes which would explain the phenomena of nature, yet always to no avail. It was me, Descartes, who broke the deadlock. It is enough to compare the suppositions of scholastics, their real qualities, substantial forms, elements and the like, whose number is almost infinite, with my supposition, i.e. with my only accepted assumption: “all bodies are composed of parts”^{7}, which in many cases is visible to the naked eye and in others can be proved by means of an infinite number of reasons.

To this statement Descartes adds that “the parts of certain kinds of bodies are of one shape rather then another”^{8}, which can be easily demonstrated to those who recognize the previous supposition. “Compare the deductions I have made from my hypothesis – about vision, salt, winds, clouds, snow, thunder, the rainbow, and so on – with what the others have derived from their hypotheses on the same topics! I hope this will be enough to convince anyone unbiased that the effects which I explain have no other causes than the ones from which I have derived them.”^{9} In these words Descartes refers directly to what he wrote about in Discourse on Method where he explains the consecutive stages of cognition, namely becoming aware of simple ideas and inferring from them basic laws of physics. From them – as from the causes – one can afterwards deduce their effects.^{10}

Thus all natural phenomena can be deduced from the basic theses of which the very primary one is the thesis of bodies being composed of

7 AT II, 200; K, 59.

8 AT II, 200; K, 59.

9 AT II, 200; K, 59.

10 See Discourse on Method, part VI.

parts, viz. that of the infinite divisibility of bodies. Expressing this thesis in different words, Descartes proponded it also as the thesis of extension constituting the essence of material bodies. The thesis was put forward as the result of cognitive processes carried out within the mind exclusively. The idea of extension is one of those simple notions which need not even be defined.^{11} But it is a natural thesis referring only to the world of natural empiricism. If it is supposed to be the departure for deduction, then a set of natural theses will be the result. From premises one can deduce conclusions of the same character that they are, hence only natural theses can logically follow from natural theses. In investigative procedure we deal with the problem of selecting statements expressing unknown causes for statements expressing known effects. In order that such a procedure be methodologically valid, the thesis that extension is the essence of bodies should be understood in the sense of the thesis of the material homogeneity of nature. And that is how Descartes understands it when he writes in the Principles of Philosophy: “There is therefore but one matter in the whole universe, and we know this by the simple fact of its being extended. All the properties which we clearly perceive in it may be reduced to the one, viz. that it can be divided or moved according to its parts, and consequently is capable of all these affections which we perceive can arise from the motion of its parts. For its partition by thought alone makes no difference to it; but all the variation in matter, or diversity in its form, depends on motion.”^{12}

The above fragment is, among others, a reflection of the mechanistic view of the universe held by Descartes. However, its significance, particularly in view of what has been said before, is much greater. After all, the mechanistic theory as such need not assume the material homogeneity of nature. Furthermore, Descartes proclaims something else: all the constitutive elements of the universe are identical in terms of their material homogeneity and it cannot be otherwise even in the very order of thought: “... the earth and heavens are formed of the same matter, and (...) even were there an infinitude of worlds, they would all be formed of this matter, from which it follows that there cannot be a plurality of worlds because we clearly perceive that the matter whose nature consists in its being an extended substance only, now occupies all the imaginable spaces where these other worlds could alone be, and we cannot find in ourselves the idea of any

11 AT II, 597; K, 65-66.

12 AT VIII-1, 52; HR I, 265.

other matter.”^{13}

As can be seen, Descartes’ idealistic epistemological standpoint leads him consequently to the materialistic ontological one. It is a standpoint of materialistic monism with reference to the world of nature, contrary to the scholastic standpoint of the plurality of being which holds that in a definite single being there are real qualities of various kinds, substantial forms, etc. Explaining the nature in terms of the plurality of being is believed to be empty rhetoric resulting from the uncritical confidence put in sensory perception. If we recognize sensory data as primary cognitive data, the only departure beyond them simply must consist in multiplying notions without the possibility of checking whether anything in extra-linguistic reality corresponds with them. In Descartes’ system the inner reality of the mind is the point from which one should proceed to the reality of nature. The ontological thesis of the material homogeneity of nature is the consequence of the idealistic epistemological thesis of the intramental character of cognition, but this is why it constitutes the methodological foundation of modern science. After Descartes one should no longer pursue science in a way other than that of explaining the material phenomena of nature by means of their causes equally material in their essence. The conviction of the existence of complete isomorphism of the structures of the mind and the universe accounted for the use of mathematics as the only language adequate for describing the realm of nature.

However, the thesis of the material homogeneity of nature, which is fundamental to science, leads on to another fundamental thesis in philosophy. In order to come to it we shall analyse the third of the indicated logical relations between cause and effect. The effect can be deduced from the cause and the cause explains the effect. The relation of proving occurs in the opposite direction: the effect proves the cause. From the thesis of the material homogeneity of nature follows the thesis of the materiality of the relation of causal determination in the realm of nature. Thus, if the effect proves the existence of the cause, it proves the existence of exclusively material causes. It is a conclusion of great philosophical importance. In the philosophical sense this conclusion should be understood as the thesis that nature does not lead outside itself, viz. rational inquiry into nature cannot lead beyond it. In other words: all efforts proceeding from the natural realm to reach the transcendental realm are invalid. After Descartes one could no longer pursue philosophy in the way characteristic of Aristotelian scholasticism, viz. trying to reach the transcendental realm by

13 AT VIII-1, 52; HR I, 265.

proceeding from the natural realm. The philosophical explanation of nature cannot lead beyond it because natural reality as such does not demand that the reason of its existence should be set forth since nature itself contains the principles explaining its character. This thesis underlies all materialistic trends in modern and contemporary philosophy.

Descartes’ great thesis on nature – the thesis that nature does not lead beyond itself in rational inquiry – was not the ultimate one in his philosophy. He assumed the existence of the transcendental realm and was convinced that one could reach it by the use of reason, only that one should not proceed from the reality of nature but from the inner reality of the mind. That is the aspect in which he formulated his second great thesis: that the mind leads outside itself demanding that the cause of its existence be given. This thesis underlines all idealistic trends in modern and contemporary philosophy. Of course, it needs separate explanation and discussion.